OpenStudy (anonymous):

The values (9.4, 11) and (11, y) are from an inverse variation. Find the missing value and round to the nearest hundredth. HELP IS APPRECIATED!

OpenStudy (amistre64):

in an inverse relation; k = xy k = 9.4 * 11 and k = 11 * y

OpenStudy (amistre64):

when does k=k?